OpenStudy (anonymous):

NUMERICAL PROBLEM: A 5gm bullet is fired from 3kg of gun. At a speed of 300 m/s. What is the speed of recoil of gun?

OpenStudy (jamesj):

Use Conservation of Momentum.

OpenStudy (ash2326):

Conserve Momentum Momentum of bullet= 5gm*300m/s=>5*10^-3*300=1.5 Kg m/s let the speed of recoil of gun be x m/s so Momentum of bullet= x *3kg=3x Kg m/s the two momentums are equal so 3x=1.5 x=0.5 m/s

OpenStudy (anonymous):

Thnx :)